A truck of 10000kg approaches a curved path with radius of 5000cm. Determine the speed limit of the angle of banking of the curves is 15°


1
Expert's answer
2021-08-29T16:55:07-0400

Given:

"R=5000\:\rm cm=50\: m"

"\theta=15^{\circ}"



The Newton's second law says:

"ma_x=N\sin\theta\\
ma_y=N\cos\theta-mg"

or

"m\frac{v^2}{R}=N\sin\theta\\
0=N\cos\theta-mg"

From these equations we get

"v=\sqrt{gR\tan\theta}=\sqrt{9.8*50*\tan 15^{\circ}}=11.5\:\rm m/s"